### v1

@@ 4. Half-Plane Property
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$. Then the corresponding $d\times d$ Hurwitz matrix is given by ${\mathrm{Hur}}_{f(t)}=\begin{bmatrix}a_{1}&a_{3}&a_{5}&\dots&\dots&\dots&0&0&0\\ a_{0}&a_{2}&a_{4}&&&&\vdots&\vdots&\vdots\\ 0&a_{1}&a_{3}&&&&\vdots&\vdots&\vdots\\ \vdots&a_{0}&a_{2}&\ddots&&&0&\vdots&\vdots\\ \vdots&0&a_{1}&&\ddots&&a_{d}&\vdots&\vdots\\ \vdots&\vdots&a_{0}&&&\ddots&a_{d-1}&0&\vdots\\ \vdots&\vdots&0&&&&a_{d-2}&a_{d}&\vdots\\ \vdots&\vdots&\vdots&&&&a_{d-3}&a_{d-1}&0\\ 0&0&0&\dots&\dots&\dots&a_{d-4}&a_{d-2}&a_{d}\end{bmatrix}$, cf., [JM12, Chap-7].
@@
Using Routh-Hurwitz criterion, cf., [BK64, Chap 4], the FAE:
@@
By expanding the determinant of ${\mathrm{Hur}}_{f(t)}$ along the elements of its last column gives $\det({\mathrm{Hur}}_{f(t)})=a_{d}\Delta_{d-1}$ where $\Delta_{k}$ represents the $k\times k$ leading principal minor of ${\mathrm{Hur}}_{f(t)}$. Since $a_{i}>0$, to demonstrate $f(t)$ of degree $d$ is Hurwitz-stable, it suffices to verify that $\Delta_{k}$ are positive for all $11\leq k\leq d-$ 1.
@@
(Liénard-Chipart) Necessary and sufficient conditions for the polynomial $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ with $a_{0}>0$ to be stable can be given in any one of the following four forms: [Gan98, pg-221]
@@
A multivariate polynomial $f(x)\in{\mathbb{R}}[x]$ is called Hurwitz-stable polynomial if all of its roots lie in the open left half plane, i.e., $f(x)\neq 0$ for all $x\in{\mathbb{R}}^{n}$ with $\mathsf{Re}(x_{i})>0$ for all $1(i)n$ or it’s a zero polynomial [COSW04]. Equivalently, a nonzero polynomial $f\in{\mathbb{R}}[x]$ is Hurwitz-stable polynomial if $f(x+tv)$ is Hurwitz-stable polynomial for any $v\in\operatorname{int}{\mathcal{K}}$ and all $x\in{\mathbb{R}}^{n}$.
@@
A polynomial $f\in{\mathbb{R}}[x]$ is said to be Hurwitz-stable polynomial over a proper convex cone ${\mathcal{K}}$ if its univariate restriction polynomial $f(x+tv)$ is Hurwitz stable for all $x\in{\mathbb{R}}^{n}$ and $v\in\operatorname{int}{\mathcal{K}}$, or it’s a zero polynomial.
@@
[BGLS01, Theorem 5.3] Any hyperbolic polynomial $f\in{\mathbb{R}}[x]^{d}_{n}$ is Hurwitz-stable over its hyperbolicity cone $\Lambda_{++}(f,e)$ containing $e$ such that $f(e)>0$.
@@
Let $f(t)=\sum_{k=0}^{n}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,n$. If $a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $11\leq k\leq d-$ 1, the following inequalities hold:
@@
    For any $21\leq k\leq d-$ 2, $a_{k}a_{k+1}\geq\frac{k+2}{k}a_{k-1}a_{k+2}\Leftrightarrow\frac{a_{k}}{a_{k-1}}\geq\frac{k+2}{k}\frac{a_{k+2}}{a_{k+1}}\Leftrightarrow\frac{a_{k+1}}{a_{k-1}}\geq\frac{k+2}{k}\frac{a_{k+2}}{a_{k}}$.
@@
Let $f(t)=\sum_{k=0}^{n}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,n$. If $a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $11\leq k\leq d-$ 1, then the following inequalities hold:
@@
    $\frac{a_{2}}{a_{0}}>\frac{a_{3}}{a_{1}}>\frac{a_{4}}{a_{2}}>\frac{a_{5}}{a_{3}}>\dots>\frac{a_{k}}{a_{k-2}}>\frac{a_{k+1}}{a_{k-1}}>\frac{a_{k+2}}{a_{k}}>\dots$.
@@
    $\frac{a_{1}}{a_{0}}>\frac{a_{3}}{a_{2}}>\frac{a_{5}}{a_{4}}>\dots\frac{a_{k-1}}{a_{k-2}}>\frac{a_{k+1}}{a_{k}}>\frac{a_{k+3}}{a_{k+2}}>\dots$
@@
    $\frac{a_{k+1}-a_{k}r}{a_{k-1}-a_{k-2}r}<\frac{a_{k+1}}{a_{k-1}}\Leftrightarrow a_{k}a_{k-1}>a_{k-2}a_{k+1}$, and  $\frac{a_{k+1}-a_{k+2}r}{a_{k-1}-a_{k}r}>\frac{a_{k+1}}{a_{k-1}}\Leftrightarrow a_{k}a_{k+1}>a_{k-1}a_{k+2}$ for any positive $r$.
@@
$(a)$ The chain of inequalities are obtained by combining Proposition 4.7 $(a)$ and $(d)$.
@@
    For any $30\leq k\leq d-$ 3, $D_{v}^{k+1}f(x)D_{v}^{k+2}f(x)\geq D_{v}^{k}f(x)D_{v}^{k+3}f(x)\Leftrightarrow\frac{D_{v}^{k+1}f(x)}{D_{v}^{k}f(x)}\geq\frac{D_{v}^{k+3}f(x)}{D_{v}^{k+2}f(x)}$.
@@
$(a)$ Since $f$ is ${\mathcal{K}}$-CLC, so the sequence $\{f(x),D_{v}f(x),D_{v}^{2}f(x),\dots,D_{v}^{d}f(x)\}$ of positive real numbers forms a log-concave sequence for any $x\in\operatorname{int}{\mathcal{K}}$ by Theorem 3.15.
$(b)$ and $(c)$ The inequalities hold for any $x\in\operatorname{int}{\mathcal{K}}$ by Proposition 4.7. ∎
@@
[BL23b, Corollary 8:13] If $f\in{\mathbb{R}}[x]$ and $g\in{\mathbb{R}}[x]$ are ${\mathcal{K}}$-CLC multivariate polynomials, the product $fg$ is ${\mathcal{K}}$-CLC polynomial
@@
Hermite-Biehler’s criterion for stability: [Lev80, Chap VII], [Kur92] Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,d$. $f(t)$ is stable if and only if $f^{e}(t)$ and $tf^{o}(t)$ have simple real interlacing zeros. Alternatively, $f(t)$ is stable if and only if all the zeros of $f^{\mathrm{even}}(t)$ and $f^{\mathrm{odd}}(t)$ are distinct, lie on the imaginary axis and alternate along it.
We use the following Hurwitz-Stability test for real polynomials based on the Interlacing Theorem and therefore, on the Boundary Crossing Theorem, cf. [BK95, Chap 1]. Note that $f(t)=f^{\mathrm{even}}(t)+f^{\mathrm{odd}}(t)$. Define the polynomial $g(t)$ of degree $d-1$ by :
@@
[BK95, Chap 1] If $f(t)$ has all its coefficients positive, $f(t)$ is stable if and only if $g(t)$ is stable
It is well known and easy to verify that polynomials of degree $1$ and $2$ with positive leading coefficients are Hurwitz-stable if and only if all their coefficients are positive. Thus the result is trivially true for $d=1$ and $d=2$. By Theorem 4.2, it’s easy to verify that for $d=3$, the $3\times 3$ leading principal minor of ${\mathrm{Hur}}_{f(t)}$ is positive if and only if $a_{1}a_{2}-a_{0}a_{3}>0$, cf. [BK95, Chap-1]. Since $a_{1}a_{2}>3a_{0}a_{3}$ by Proposition 4.7, so $f(t)$ is Hurwitz-stable.
@@
Let $f(t)=\sum_{k=0}^{4}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,4$. If $a_{k}^{2}>\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $11\leq k\leq d-$ 1, then $f(t)$ is Hurwitz-stable polynomial.
@@
For $d=4$, by Theorem 4.13, $f(t)=\sum_{k=0}^{4}a_{k}t^{k}$ is Hurwitz-stable if and only if $g(t):=f(t)-\mu tf^{\mathrm{odd}}(t)=a_{0}+a_{1}t+(a_{2}-\frac{a_{4}}{a_{3}}a_{1})t^{2}+a_{3}t^{3}$ is Hurwitz-stable, where $\mu=\frac{a_{4}}{a_{3}}$. Note that $g(t)$ is a cubic univariate polynomial with positive coefficients since $a_{2}a_{3}>a_{1}a_{4}$ by Proposition 4.7(a). Thus, cubic polynomial $g(t)$ is Hurwitz-stable if and only if $00<\frac{a_{0}}{a_{2}-\frac{a_{4}}{a_{3}}a_{1}}<\frac{a_{1}}{a_{3}}\Leftrightarrow a_{1}a_{2}a_{3}-a_{1}^{2}a_{4}-a_{0}a_{3}^{2}>0\Leftrightarrow a_{1}(a_{2}a_{3}-2a_{1}a_{4})+a_{1}^{2}a_{4}-a_{0}a_{3}^{2}>$ 0. Furthermore,
$a_{1}(a_{2}a_{3}-2a_{1}a_{4})+2a_{0}a_{2}a_{4}-a_{0}a_{3}^{2}>0$
$\Leftrightarrow a_{1}(a_{2}a_{3}-2a_{1}a_{4})>a_{0}(a_{3}^{2}-2a_{2}a_{4})$
$\Leftrightarrow\frac{a_{1}}{a_{0}}>\frac{a_{3}}{a_{2}}\frac{a_{3}-2\frac{a_{2}}{a_{3}}a_{4}}{a_{3}-2\frac{a_{1}}{a_{2}}a_{4}}$
Note that the denominator and numerator in the rhs are all positive and $\frac{a_{3}-2\frac{a_{2}}{a_{3}}a_{4}}{a_{3}-2\frac{a_{1}}{a_{2}}a_{4}}<1$ since $\frac{a_{2}}{a_{3}}>\frac{a_{1}}{a_{2}}$. Thus, the strict inequality $\frac{a_{1}}{a_{0}}>\frac{a_{3}}{a_{2}}>\frac{a_{3}}{a_{2}}\frac{a_{3}-2\frac{a_{2}}{a_{3}}a_{4}}{a_{3}-2\frac{a_{1}}{a_{2}}a_{4}}$ holds because $a_{1}a_{2}>a_{0}a_{3}$. On the other hand, since $a_{1}^{2}>2a_{0}a_{2}$, so $a_{1}(a_{2}a_{3}-2a_{1}a_{4})+a_{1}^{2}a_{4}-a_{0}a_{3}^{2}>a_{1}(a_{2}a_{3}-2a_{1}a_{4})+2a_{0}a_{2}a_{4}-a_{0}a_{3}^{2}>0$. Thus, $g(t)$ is Hurwitz-stable. Consequently, $f(t)$ is Hurwitz-stable. ∎
Therefore, we can conclude the following result.

### v2

@@ 4. Half-Plane Property
Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$. Then the corresponding $d\times d$ Hurwitz matrix is given by ${\mathrm{Hur}}_{f(t)}=\begin{bmatrix}a_{1}&a_{3}&a_{5}&\dots&\dots&\dots&0&0&0\\ a_{0}&a_{2}&a_{4}&&&&\vdots&\vdots&\vdots\\ 0&a_{1}&a_{3}&&&&\vdots&\vdots&\vdots\\ \vdots&a_{0}&a_{2}&\ddots&&&0&\vdots&\vdots\\ \vdots&0&a_{1}&&\ddots&&a_{d}&\vdots&\vdots\\ \vdots&\vdots&a_{0}&&&\ddots&a_{d-1}&0&\vdots\\ \vdots&\vdots&0&&&&a_{d-2}&a_{d}&\vdots\\ \vdots&\vdots&\vdots&&&&a_{d-3}&a_{d-1}&0\\ 0&0&0&\dots&\dots&\dots&a_{d-4}&a_{d-2}&a_{d}\end{bmatrix}$, cf., [stabilityHurwitz, Chap-7].
@@
Using Routh-Hurwitz criterion, cf., [Bellmanselected, Chap 4], the FAE:
@@
By expanding the determinant of ${\mathrm{Hur}}_{f(t)}$ along the elements of its last column gives $\det({\mathrm{Hur}}_{f(t)})=a_{d}\Delta_{d-1}$ where $\Delta_{k}$ represents the $k\times k$ leading principal minor of ${\mathrm{Hur}}_{f(t)}$. Since $a_{i}>0$, to demonstrate $f(t)$ of degree $d$ is Hurwitz-stable, it suffices to verify that $\Delta_{k}$ are positive for all $1\leq k\leq d-1$.
@@
(Liénard-Chipart) Necessary and sufficient conditions for the polynomial $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ with $a_{0}>0$ to be stable can be given in any one of the following four forms: [gantmacher, pg-221]
@@
A polynomial $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$ is called (Hurwitz-)stable if $f(z)\neq 0$ for all $z=(z_{1},\dots,z_{n})\in{\mathbb{C}}^{n}$ with $\mathsf{Re}(z_{i})>0$ for all $1\leq i\leq n$. Equivalently, all zeros of $f$ lie in the closed left half-plane in each coordinate. By convention, we also regard the zero polynomial as stable.
Equivalently, if $f\in{\mathbb{R}}[x_{1},\dots,x_{n}]$ is homogeneous and nonzero, then $f$ is Hurwitz-stable if and only if, for all $x,v\in({\mathbb{R}}_{>0})^{n}$, the univariate polynomial
$t\ \longmapsto\ f(x+tv)$
is a Hurwitz-stable polynomial in $t$, that is, all of its zeros have negative real part; see, e.g., [Gregor, Lemma 6], [Oxley, Prop. 5.2].
@@
Let ${\mathcal{K}}\subset{\mathbb{R}}^{n}$ be a proper convex cone with nonempty interior $\operatorname{int}{\mathcal{K}}$. A polynomial $f\in{\mathbb{R}}_{n}^{d}[x]$ is said to be Hurwitz-stable over $\mathcal{K}$} if either $f\equiv 0$, or else for every $x,v\in\operatorname{int}{\mathcal{K}}$ the univariate polynomial
$t\ \longmapsto\ f(x+tv)$
is Hurwitz-stable in the univariate sense, i.e., all of its zeros have negative real part.
@@
[Gulerhyperbolicconvex, Theorem 5.3] Any hyperbolic polynomial $f\in{\mathbb{R}}[x]^{d}_{n}$ is Hurwitz-stable over its hyperbolicity cone $\Lambda_{++}(f,e)$ containing $e$ such that $f(e)>0$.
Here we derive Newton-type inequalities for directional derivatives, which we then feed into stability criteria.
@@
Let $f(t)=\sum_{k=0}^{n}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,n$. If $a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $1\leq k\leq d-1$, the following inequalities hold:
@@
    For any $1\leq k\leq d-2$, $a_{k}a_{k+1}\geq\frac{k+2}{k}a_{k-1}a_{k+2}\Leftrightarrow\frac{a_{k}}{a_{k-1}}\geq\frac{k+2}{k}\frac{a_{k+2}}{a_{k+1}}\Leftrightarrow\frac{a_{k+1}}{a_{k-1}}\geq\frac{k+2}{k}\frac{a_{k+2}}{a_{k}}$.
@@
Let $f(t)=\sum_{k=0}^{n}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,n$. If $a_{k}^{2}\geq\frac{k+1}{k}a_{k-1}a_{k+1}$ for all $1\leq k\leq d-1$, then the following inequalities hold:
@@
    Monotonic chains of adjacent ratios: $\frac{a_{2}}{a_{0}}>\frac{a_{3}}{a_{1}}>\frac{a_{4}}{a_{2}}>\frac{a_{5}}{a_{3}}>\dots>\frac{a_{k}}{a_{k-2}}>\frac{a_{k+1}}{a_{k-1}}>\frac{a_{k+2}}{a_{k}}>\dots$.
@@
    Separated-index ratio chains: $\frac{a_{1}}{a_{0}}>\frac{a_{3}}{a_{2}}>\frac{a_{5}}{a_{4}}>\dots\frac{a_{k-1}}{a_{k-2}}>\frac{a_{k+1}}{a_{k}}>\frac{a_{k+3}}{a_{k+2}}>\dots$
@@
    Cross-ratio type inequalities: $\frac{a_{k+1}-a_{k}r}{a_{k-1}-a_{k-2}r}<\frac{a_{k+1}}{a_{k-1}}\Leftrightarrow a_{k}a_{k-1}>a_{k-2}a_{k+1}$, and $\frac{a_{k+1}-a_{k+2}r}{a_{k-1}-a_{k}r}>\frac{a_{k+1}}{a_{k-1}}\Leftrightarrow a_{k}a_{k+1}>a_{k-1}a_{k+2}$ for any positive $r$.
@@
$(a)$ The chain of inequalities is obtained by combining Proposition˜4.7 $(a)$ and $(d)$.
@@
Note that strong log-concavity $\Rightarrow$ very rigid ordering for coefficient ratios and related rational expressions. We translate abstract log-concavity into very concrete, ordered ratio inequalities that can be plugged directly into determinant conditions for Hurwitz matrices.
@@
    For any $0\leq k\leq d-3$, $D_{v}^{k+1}f(x)D_{v}^{k+2}f(x)\geq D_{v}^{k}f(x)D_{v}^{k+3}f(x)\Leftrightarrow\frac{D_{v}^{k+1}f(x)}{D_{v}^{k}f(x)}\geq\frac{D_{v}^{k+3}f(x)}{D_{v}^{k+2}f(x)}$.
@@
$(a)$ Since $f$ is $\mathcal{K}$}-CLC, so the sequence $\{f(x),D_{v}f(x),D_{v}^{2}f(x),\dots,D_{v}^{d}f(x)\}$ of positive real numbers forms a log-concave sequence for any $x\in\operatorname{int}{\mathcal{K}}$ by Theorem˜3.16.
$(b)$ and $(c)$ The inequalities hold for any $x\in\operatorname{int}{\mathcal{K}}$ by Proposition˜4.7. ∎
@@
[Hereditary, Corollary 8:13] If $f\in{\mathbb{R}}[x]$ and $g\in{\mathbb{R}}[x]$ are $\mathcal{K}$}-CLC multivariate polynomials, the product $fg$ is $\mathcal{K}$}-CLC polynomial
@@
Hermite-Biehler’s criterion for stability: [Levindistribution, Chap VII], [Kurtzsufficient] Let $f(t)=\sum_{k=0}^{d}a_{k}t^{k}$ be a univariate polynomial with positive $a_{k}$ for $k=0,\dots,d$. $f(t)$ is stable if and only if $f^{e}(t)$ and $tf^{o}(t)$ have simple real interlacing zeros. Alternatively, $f(t)$ is stable if and only if all the zeros of $f^{\mathrm{even}}(t)$ and $f^{\mathrm{odd}}(t)$ are distinct, lie on the imaginary axis and alternate along it.
We use the following Hurwitz-Stability test for real polynomials based on the Interlacing Theorem and therefore, on the Boundary Crossing Theorem, cf. [Bhattacharyyarobust, Chap 1]. Note that $f(t)=f^{\mathrm{even}}(t)+f^{\mathrm{odd}}(t)$. Define the polynomial $g(t)$ of degree d1d-1 by :
@@
[Bhattacharyyarobust, Chap 1] If $f(t)$ has all its coefficients positive, $f(t)$ is stable if and only if $g(t)$ is stable
It is well known and easy to verify that polynomials of degree $1$ and $2$ with positive leading coefficients are Hurwitz-stable if and only if all their coefficients are positive. Thus the result is trivially true for $d=1$ and $d=2$. By Theorem˜4.2, it’s easy to verify that for $d=3$, the $3\times 3$ leading principal minor of ${\mathrm{Hur}}_{f(t)}$ is positive if and only if $a_{1}a_{2}-a_{0}a_{3}>0$, cf. [Bhattacharyyarobust, Chap-1]. Since $a_{1}a_{2}>3a_{0}a_{3}$ by Proposition˜4.7, so $f(t)$ is Hurwitz-stable.
@@
Let $f(t)=\sum_{k=0}^{4}a_{k}t^{k}$ be a quartic polynomial with $a_{k}>0$ for $k=0,\dots,4$. Suppose that
$a_{k}^{2}\;>\;\frac{k+1}{k}\,a_{k-1}a_{k+1}\qquad\text{for }k=1,2,3.$
Then $f(t)$ is Hurwitz–stable.
@@
By Theorem˜4.13, $f(t)=\sum_{k=0}^{4}a_{k}t^{k}$ is Hurwitz–stable if and only if
$g(t)\;:=\;f(t)-\mu tf^{\mathrm{odd}}(t)\;=\;a_{0}+a_{1}t+\Bigl(a_{2}-\frac{a_{4}}{a_{3}}a_{1}\Bigr)t^{2}+a_{3}t^{3}$
is Hurwitz–stable, where $\mu=a_{4}/a_{3}$. Set
$b_{0}:=a_{0},\quad b_{1}:=a_{1},\quad b_{2}:=a_{2}-\frac{a_{4}}{a_{3}}a_{1},\quad b_{3}:=a_{3},$
so $g(t)=b_{0}+b_{1}t+b_{2}t^{2}+b_{3}t^{3}$.
First, $b_{2}>0$ by Proposition˜4.7(a), which gives
$a_{2}a_{3}\;>\;a_{1}a_{4}\quad\Longrightarrow\quad a_{2}-\frac{a_{4}}{a_{3}}a_{1}>0.$
Hence $b_{0},b_{1},b_{2},b_{3}$ are all positive.
For a cubic with positive coefficients $b_{0},b_{1},b_{2},b_{3}>0$, the Routh–Hurwitz criterion says that Hurwitz stability is equivalent to
$b_{1}b_{2}\;>\;b_{0}b_{3}\quad\Longleftrightarrow\quad a_{1}\Bigl(a_{2}-\frac{a_{4}}{a_{3}}a_{1}\Bigr)\;>\;a_{0}a_{3}.$
Equivalently,
$a_{1}a_{2}a_{3}-a_{1}^{2}a_{4}-a_{0}a_{3}^{2}\;>\;0.$
Thus it remains to show
(1) $a_{1}a_{2}a_{3}\;>\;a_{1}^{2}a_{4}+a_{0}a_{3}^{2}.$
Define the dimensionless quantities
$r_{1}:=\frac{a_{1}^{2}a_{4}}{a_{1}a_{2}a_{3}}=\frac{a_{1}a_{4}}{a_{2}a_{3}},\qquad r_{2}:=\frac{a_{0}a_{3}^{2}}{a_{1}a_{2}a_{3}}=\frac{a_{0}a_{3}}{a_{1}a_{2}}.$
Then (1) is equivalent to
$1-r_{1}-r_{2}\;>\;0.$
From the assumed inequalities
$a_{1}^{2}>2a_{0}a_{2},\qquad a_{2}^{2}>\frac{3}{2}a_{1}a_{3},\qquad a_{3}^{2}>\frac{4}{3}a_{2}a_{4},$
we obtain the following bounds.
1. For $r_{2}$:
$\frac{a_{0}}{a_{1}}<\frac{a_{1}}{2a_{2}}\quad\Rightarrow\quad r_{2}=\frac{a_{0}}{a_{1}}\cdot\frac{a_{3}}{a_{2}}<\frac{a_{1}}{2a_{2}}\cdot\frac{a_{3}}{a_{2}}=\frac{a_{1}a_{3}}{2a_{2}^{2}}.$
Using $a_{2}^{2}>\frac{3}{2}a_{1}a_{3}$ gives
$\frac{a_{1}a_{3}}{a_{2}^{2}}<\frac{2}{3}\quad\Rightarrow\quad r_{2}<\frac{1}{2}\cdot\frac{2}{3}=\frac{1}{3}.$
2. For $r_{1}$:
$r_{1}=\frac{a_{1}}{a_{2}}\cdot\frac{a_{4}}{a_{3}}.$
From $a_{2}^{2}>\frac{3}{2}a_{1}a_{3}$, $\frac{a_{1}}{a_{2}}<\frac{2}{3}\cdot\frac{a_{2}}{a_{3}}$, and from $a_{3}^{2}>\frac{4}{3}a_{2}a_{4}$, $\frac{a_{4}}{a_{3}}<\frac{3}{4}\cdot\frac{a_{3}}{a_{2}}$. Multiplying these estimates gives
$r_{1}<\frac{2}{3}\cdot\frac{3}{4}\cdot\frac{a_{2}}{a_{3}}\cdot\frac{a_{3}}{a_{2}}=\frac{1}{2}.$
Combining the two bounds, we get
$r_{1}+r_{2}<\frac{1}{2}+\frac{1}{3}=\frac{5}{6}<1,$
so $1-r_{1}-r_{2}>0$, which is exactly (1). Hence $g$ is Hurwitz–stable, and therefore $f$ is Hurwitz–stable by Theorem˜4.13. ∎